how we can find hybridization for AsO3^3-, CO3^2-, ClO3-, BO3^3-

Grade:12th pass

1 Answers

sadaf sagheer
5 years ago

for AsO3^3-,firslty find total valence electrons= 5+ 6(3)-(-3)=26.then 26/8, divide remainder by 2, and then add quotient and remainder/2

i.e 3+1= 4 hence the hybridisation is sp3.for All has same methods.